# Learning to Manipulate under Limited Information

### v2

@@ Abstract.
By classic results in social choice theory, any reasonable preferential voting method sometimes gives individuals an incentive to report an insincere preference. The extent to which different voting methods are more or less resistant to such strategic manipulation has become a key consideration for comparing voting methods. Here we measure resistance to manipulation by whether neural networks of varying sizes can learn to profitably manipulate a given voting method in expectation, given different types of limited information about how other voters will vote. We trained over 70,000 neural networks of 26 sizes to manipulate against 8 different voting methods, under 6 types of limited information, in committee-sized elections with 5–21 voters and 3–6 candidates. We find that some voting methods, such as Borda, are highly manipulable by networks with limited information, while others, such as Instant Runoff, are not, despite being quite profitably manipulated by an ideal manipulator with full information. For the two probability models for elections that we use, the overall least manipulable of the 8 methods we study are Condorcet methods, namely Minimax and Split Cycle.
Key words and phrases:
computational social choice, strategic voting, machine learning
1. Introduction
@@
A fundamental problem in multi-agent decision making is that of aggregating heterogeneous preferences Conitzer (2010). Voting theory provides many possible methods of preference aggregation with different benefits and costs Zwicker (2016). However, no reasonable preferential voting method escapes the problem of manipulability. As shown by classic results such as the Gibbard-Satterthwaite theorem Gibbard (1973); Satterthwaite (1973) and its generalizations Duggan and Schwartz (2000); Taylor (2005), for any such voting method, there is some preference profile in which some voter has an incentive to report an insincere preference in order to obtain a result that is preferable, according to their sincere preference, to the result they would obtain if they were to submit their sincere preference. Thus, sincere voting is not a Nash equilibrium of the game derived from this preference profile, where the players are the voters and the actions are the possible preference rankings to report.
@@
However, the mere existence of a preference profile in which a voter has an incentive to misreport their preferences tells us little about the frequency with which a voter will have such an incentive or the difficulty of recognizing that such an incentive exists—under either full information or limited information about the profile. Thus, a manipulable voting method might be relatively resistant to manipulation, either because the frequency just cited is low, the difficulty just cited is high, or a mixture of both factors. Such resistance to manipulation has been considered an important criterion for comparing voting rules Merrill (1988); Green-Armytage et al. (2016).
@@
As for the difficulty of manipuation, there is now a large literature in computational social choice on the worst-case complexity of computing whether there is a strategic ranking that will elect a desired candidate. Faliszewski and Procaccia Faliszewski and Procaccia (2010) call this literature “AI’s war on manipulation.” A series of hardness results have been proved since Bartholdi et al. (1989) and Conitzer et al. (2007). There is evidence for and against the view that high worst-case computational complexity is a barrier to manipulation Walsh (2011), and the situation with average-case manipulability might be quite different Mossel and Rácz (2013).
@@
In this paper, we take a different approach than previous work, measuring resistance to manipulation by whether neural networks of varying sizes can learn to profitably manipulate a given voting method in expectation, given different types of information about how other voters will vote. Like the classic results on manipulation, we focus on the case of a single manipulating voter. A single voter can almost never affect the outcome of a large election with thousands of voters or more by changing their vote, so rational manipulation by a single voter is most relevant for small elections in committees, boards, etc. The classic results on manipulation also in effect assume that the manipulator knows exactly how all other voters will vote, which is unrealistic in many voting contexts. By contrast, we train neural networks to manipulate on the basis of different types of limited information. [^1] All code and Supplementary Material is available at https://github.com/epacuit/ltm.
1.1. Related work
1.1.1. Manipulation under Limited Information
@@
A number of previous papers study whether a voter can successfully manipulate in an election under limited information about how other voters will vote Myerson and Weber (1993); Conitzer et al. (2011); Reijngoud and Endriss (2012); Meir et al. (2014); Endriss et al. (2016); Lang (2020); Veselova (2023), including with the use of heuristics Chopra et al. (2004); Laslier (2009); Meir (2018); Fairstein et al. (2019). The information of the manipulator is typically represented by a set of preference profiles, all of which agree on (i) the manipulator’s own preferences and (ii) some other partial information (e.g., all profiles in the manipulator’s information set agree on who will win the election if the manipulator votes sincerely, or all profiles in the set are such that for each of the other voters $i$, $i$’s ranking in the profile extends some known partial order over the candidates that the manipulator attributes to $i$), perhaps supplemented with a probability measure over such profiles Lu et al. (2012). By contrast, in this paper we represent limited information by the inputs to a neural network, as described below.
1.1.2. Machine Learning and Voting Theory
@@
Several previous works have applied machine learning to problems in voting theory, though not in the way we do here. Anil and Bao (2021) and Burka et al. (2022) study the learnability of various voting methods, Kang et al. (2023) studies learning how to explain election results for various voting methods, and Armstrong and Larson (2019) and Firebanks-Quevedo (2020) use machine learning to create new voting methods satisfying desiderata, but none of these papers discuss learning to manipulate as a voter. Manipulation is studied in Airiau et al. (2017), but only in the context of iterative voting, [^2] In iterative voting, after all voters submit rankings and a tentative winner is announced, the voters are allowed to sequentially change their rankings, with a new tentative winner announced after each change, until an equilibrium is reached. whereas here we focus on learning to manipulate in traditional elections, where the final winner is immediately computed after all voters submit their rankings.
1.1.3. Machine Learnability as a Metric of Task Difficulty
@@
Sufficiently large neural networks have been shown to be able to learn arbitrarily complex functions, including fitting to random data Zhang et al. (2017). If the model is not large enough to fully memorize the training data, however, learning the training data requires generalization. In the fields of reinforcement learning and natural language processing, it is commonly held that more complex problems may require larger and more complex networks. Indeed, previous work has shown that model performance grows as the number of learnable parameters increases Hestness et al. (2017); Golubeva et al. (2021). In this paper, we use required model size as a proxy for task difficulty. Not only has learnability by neural networks been taken to be suggestive of human learnability Steinert-Threlkeld and Szymanik (2020), but also, in practice, humans may use a neural network to help them manipulate an election under limited information.
2. Preliminaries
@@
Given a set $V$ of voters and a set $X$ of candidates, a preference profile for $(V,X)$ is a function $\mathbf{P}$ assigning to each $i\in V$ a linear order $\mathbf{P}_{i}$ of $X$. Where $\mathbf{Q}$ is a preference profile for $(V,X)$, $i\not\in V$, and $\mathbf{P}_{i}$ is a linear order of $X$, we write $(\mathbf{P}_{i},\mathbf{Q})$ for the preference profile that assigns to $i$ the linear order $\mathbf{P}_{i}$ and assigns to each $j\in V$ the linear order $\mathbf{Q}_{j}$. For a profile $\mathbf{P}$, let $\mathbf{P}_{-i}$ be the profile restricted to all voters except $i$, so we may write $\mathbf{P}=(\mathbf{P}_{i},\mathbf{P}_{-i})$.
@@
A utility profile for $(V,X)$ is a function $\mathbf{U}$ assigning to each $i\in V$ a utility function $\mathbf{U}_{i}:X\to\mathbb{R}$, where we assume that ${\mathbf{U}_{i}(x)\neq\mathbf{U}_{i}(y)}$ whenever $x\neq y$ [^3] The probability models for generating profiles that we use make a tie in the utilities of distinct candidates a measure zero event. Given such a utility profile $\mathbf{U}$, its induced preference profile $\mathbf{P}(\mathbf{U})$ assigns to each $i\in V$ the linear order $\succ_{i}$ defined by
$x\succ_{i}y$ iff $\mathbf{U}_{i}(x)>\mathbf{U}_{i}(y)$.
@@
We say $\mathbf{P}_{i}^{\prime}$ is optimal if the left-hand side of (1) is maximized for $\mathbf{P}^{\prime}_{i}$ among all possible linear orders of $X$. We assume, as in standard decision theory Kreps (1988), that our manipulating agent aims to maximize expected utility and hence aims to submit an optimal ranking.
A voting method $F$ is manipulable at $\mathbf{U}$ by $i$ if there is some profitable manipulation of $F$ at $\mathbf{U}$ by $i$; and $F$ is manipulable if there is some utility profile $\mathbf{U}$ and voter $i$ such that $F$ is manipulable at $\mathbf{U}$ by $i$. This notion of manipulability of $F$ coincides with the notion of manipulability of $F$ in Gibbard (1977) when we regard $F$ as a probabilistic voting method that assigns to each profile $\mathbf{P}$ the lottery $F_{\ell}(\mathbf{P})$.
2.1. Voting Methods
@@
Instant Runoff (IRV): if more than half the voters rank the same candidate $A$ in first place, then $A$ wins; otherwise, $A$ is an IRV winner if $A$ is an IRV winner in the profile $\mathbf{P}^{\prime}$ obtained from $\mathbf{P}$ by removing all candidates who received the fewest first-place votes in $\mathbf{P}$ (unless all candidates have the same number of first-place votes, in which case all of them win).
Instant Runoff with parallel-universe tiebreaking (IRV-PUT): if more than half of the voters rank the same candidate $A$ in first place, then $A$ wins; otherwise a candidate $A$ is an IRV-PUT winner if for one of the candidates $B$ who received the fewest first-place votes in $\mathbf{P}$, $A$ is the IRV-PUT winner in the profile $\mathbf{P}_{-B}$ obtained from $\mathbf{P}$ by removing $B$.
Minimax: where the margin of $A$ vs. $B$ in $\mathbf{P}$ is the number of voters who rank $A$ above $B$ minus the number who rank $B$ above $A$ in $\mathbf{P}$, the winners are those $A$ who minimize the quantity $\mathrm{max}\{margin_{\mathbf{P}}(B,A)\mid B\in X\}$.
@@
Plurality, Borda, and IRV are perhaps the most famous of preferential voting methods. Plurality has been used for many centuries, and Borda and IRV date back to at least the 18th century. The definition of IRV above (with simultaneous elimination of all candidates with the fewest first-place votes) matches that of Taylor and Pacelli (2008), while the PUT version is popular in computational social choice (see Wang et al. (2019)).
Minimax Simpson (1969); Kramer (1977) is one of the most well known of the Condorcet-consistent voting methods. A voting method is Condorcet-consistent if whenever there is some candidate $C$ who beats every other candidate $B$ head-to-head (i.e., $C$ has a positive margin over each distinct $B$), then this candidate—called the Condorcet winner—is the unique winner of the election. Plurality, Borda, IRV, and IRV-PUT all violate Condorcet consistency.
The Nanson [^4] There are two versions of Nanson, one that removes all candidates with below average Borda score (Strict Nanson) and one that removes all candidates with less than or equal to average Borda score (Weak Nanson) unless doing so eliminates everyone. Apparently Nanson himself had in mind Weak Nanson Niou (1987), but most of the literature in computational social choice focuses on Strict Nanson (see Brandt et al. (2016)), so we do as well. method Nanson (1882) is also Condorcet-consistent and has previously been studied in connection with strategic voting. In Narodytska et al. (2011), it is shown that the problem of manipulating Nanson (and the related Baldwin method) is NP-hard when the number of candidates is allowed to increase.
Finally, we include the recently proposed Split Cycle voting method Holliday and Pacuit (2023a), whose manipulability has been studied in Durand (2023), as well as one of its refinements, [^5] A voting method $F$ refines a method $G$ if for any profile $\mathbf{P}$, $F(\mathbf{P})\subseteq G(\mathbf{P})$. Other refinements of Split Cycle include Beat Path Schulze (2011) and Ranked Pairs Tideman (1987). the Stable Voting method Holliday and Pacuit (2023b). These methods satisfy not only Condorcet consistency but also the stronger property—violated by Minimax—of Smith consistency, meaning that their winners always belong to the Smith set, the smallest set of candidates such that every candidate inside the set beats every candidate outside the set head-to-head. No previous work has studied the manipulability of Stable Voting, so studying this method tests if our approach can predict the manipulability of a method as measured in other ways in the future.
3. Learning to Manipulate
@@
3.1. Implementation Details
We optimize weights $\theta$ of an MLP $f$ whose input $x$ consists of a utility function $\mathbf{U}_{v}$ for our manipulating voter $v$, as well as some limited information $I$ about the full utility profile $\mathbf{U}$ (see Section 3.1.2). We apply a softmax to the output of the MLP to generate a probability distribution over all possible actions of $v$, namely the $m!$ possible rankings, labeled 0 through $m!-1$, that $v$ can submit, as in (2):
@@ 3.1. Implementation Details
Note that the manipulator is allowed to submit the sincere ranking given by their utility function; they are not forced to manipulate.
3.1.1. Probability models for profiles
To generate utility profiles for our experiments described below, we first used a standard uniform utility model (see, e.g., (Merrill, 1988, p. 16)): for each voter independently, the utility of each candidate for that voter is drawn independently from the uniform distribution on the $[0,1]$ interval. We then also used a 2D spatial model: each candidate and each voter is independently placed in $\mathbb{R}^{2}$ according to the multivariate normal distribution (as in Merrill (1988)) with no correlation between the two dimensions; then the utility of a candidate for a voter is the square of the Euclidean distance between the candidate and the voter (using the quadratic proximity utility function as in (Samuel Merrill and Grofman, 1999, p. 21)).
These utility profiles can then be parameterized as 2D matrices, $U\in\mathbb{R}^{n\times m}$, where $n$ is the number of voters, $m$ is the number of candidates, and $U_{\left[i,c\right]}\in\left[0,1\right]$. To select the utility function for a given voter $i$ is to select the row $U_{i}$.
3.1.2. Choices of Limited Information
We experimented with providing different types of input to the MLP, including the following types that are often taken to be natural forms of polling information in voting theory (see, e.g., (Reijngoud and Endriss, 2012, § 2.2), Veselova (2023)):
@@ 3.1.2. Choices of Limited Information
These additional inputs are flattened and concatenated to the manipulator’s utility function before being used as input to the MLP.
Note, crucially, that the full preference profile $\mathbf{P}$ is not uniquely determined by any of the types of limited information above. If the manipulating voter $v$ had full knowledge of $\mathbf{P}$, they could simply compute which of the $|X|!$ linear orders would be optimal to submit given $\mathbf{P}_{-i}$. But if $v$ has only, e.g., the majority matrix of $\mathbf{P}$, then the set of profiles realizing that majority matrix can be enormous. For 6 candidates and 20 voters in addition to $v$, there are approximately $7.48\times 10^{38}$ profiles up to permuting the voters (which does not affect the choice of winners according to the anonymous voting rules in Section 2.1). [^6] This is the number of anonymous profiles for 6 candidates and 20 voters. For $m$ candidates and $n$ voters, there are $(n+m!-1)!/((m!-1)!n!)$ anonymous profiles. For 6 candidates, there are $3563^{6\choose 2}=265,$ 356 majority matrices. Thus, on average there are about $7.48\times 10^{38}/265,356\approx 2.82\times 10^{33}$ profiles that could realize a given majority matrix. It is therefore infeasible to check for every profile that could realize a given majority matrix, and every ranking that $v$ could submit, the profitability of submitting that ranking in that profile. Instead, we will train an MLP to input a majority matrix and output an optimal ranking to submit without such infeasible brute force checking.
@@ 3.1.3. Labeling
  • optimizing labeling: all optimal rankings to submit are labeled by $1$, and all other rankings are labeled by 00.777We also duplicated our experiments with a satisficing labeling (if there are profitable manipulations, they are labeled by $1$, and all others rankings are labeled by 00; otherwise all rankings that do at least as well as the sincere ranking are labeled by $1$ and all others by 00), but the MLPs trained with this labeling had qualitatively indistinguishable results from those trained with the optimizing labeling.
The output of our MLPs is a distribution over all $m!$ rankings given some information $x$ about the current utility profile. It is equally valid for our agent to choose any of the positively-labeled rankings. We treat the binary labelings as a mask over the rankings and reduce the distribution $\pi(x)$ to two values: the probability of choosing a positively-labeled ranking or not. We compute the final loss as the mean-squared error between this reduced distribution and the distribution assigning probability 1 to choosing a positively-labeled ranking and 0 to choosing a non-positively-labeled ranking.
3.2. Evaluation
@@
  • argmax: select the ranking with the largest probability in the output of the MLP.
@@
The normalization in (3) is the standard normalization for relative utilitarianism Dhillon and Mertens (1999), which we use to compare utility differences across profiles. Indeed, (3) is equivalent to taking the difference in the Kaplan-normalized (d’Aspremont and Gevers, 2002, p. 470) expected utilities of submitting $\mathbf{P}^{\prime}$ and of submitting $\mathbf{P}$, where $\hat{u}=\mathrm{max}(\{\mathbf{U}_{i}(x)\mid x\in X\})$ and $\check{u}=\mathrm{min}(\{\mathbf{U}_{i}(x)\mid x\in X\})$:
@@
For a given trained MLP, we sample utility profiles according to one of our probability models and compute the average profitability of the MLP’s submitted rankings. We take the average profitability that is achievable against a given voting method as a measure of a voter’s incentive to manipulate against that voting method. This is a more revealing measure than the frequency with which the trained MLP manipulates against a voting method, since many of the MLP’s manipulations may be no better than sincere voting.
@@
For the number of profiles to sample, we continued sampling until the estimated standard error of the mean (Robert and Casella, 2004, § 3.2) for profitability fell below 5e-4, resulting in small error bars (see Figure 3). [^8] The number of profiles sampled for evaluation of an MLP was at least 4,096 with an average of approximately 45,000 and a standard deviation of approximately 35,000.
3.2.1. Baselines

### v3

@@ Abstract
By classic results in social choice theory, any reasonable preferential voting method sometimes gives individuals an incentive to report an insincere preference. The extent to which different voting methods are more or less resistant to such strategic manipulation has become a key consideration for comparing voting methods. Here we measure resistance to manipulation by whether neural networks of various sizes can learn to profitably manipulate a given voting method in expectation, given different types of limited information about how other voters will vote. We trained over 100,000 neural networks of 26 sizes to manipulate against 8 different voting methods, under 6 types of limited information, in committee-sized elections with 5–21 voters and 3–6 candidates. We find that some voting methods, such as Borda, are highly manipulable by networks with limited information, while others, such as Instant Runoff, are not, despite being quite profitably manipulated by an ideal manipulator with full information. For the three probability models for elections that we use, the overall least manipulable of the 8 methods we study are Condorcet methods, namely Minimax and Split Cycle.
Codehttps://github.com/epacuit/ltm
Introduction
@@
A fundamental problem in multi-agent decision making is that of aggregating heterogeneous preferences (Conitzer 2010). Voting theory provides many possible methods of preference aggregation with different benefits and costs (Zwicker 2016). However, no reasonable preferential voting method escapes the problem of manipulability. As shown by classic results such as the Gibbard-Satterthwaite theorem (Gibbard 1973; Satterthwaite 1973) and its generalizations (Duggan and Schwartz 2000; Taylor 2005), for any such voting method, there is some preference profile in which some voter has an incentive to report an insincere preference in order to obtain a result that is preferable, according to their sincere preference, to the result they would obtain if they were to submit their sincere preference. Thus, sincere voting is not a Nash equilibrium of the game derived from this preference profile, where the players are the voters and the actions are the possible preference rankings to report.
@@
However, the mere existence of a preference profile in which a voter has an incentive to misreport their preferences tells us little about the frequency with which a voter will have such an incentive or the difficulty of recognizing that such an incentive exists—under either full information or limited information about the profile. Thus, a manipulable voting method might be relatively resistant to manipulation, either because the frequency just cited is low, the difficulty just cited is high, or a mixture of both factors. Such resistance to manipulation has been considered an important criterion for comparing voting rules (Merrill 1988; Green-Armytage, Tideman, and Cosman 2016).
@@
As for the difficulty of manipuation, there is now a large literature in computational social choice on the worst-case complexity of computing whether there is a strategic ranking that will elect a desired candidate. Faliszewski and Procaccia (2010) call this “AI’s war on manipulation.” A series of hardness results have been proved since Bartholdi, Tovey, and Trick 1989 and Conitzer, Sandholm, and Lang 2007. There is evidence for and against the view that high worst-case computational complexity is a barrier to manipulation (Walsh 2011), and the situation with average-case manipulability might be quite different (Mossel and Rácz 2013).
@@
In this paper, we take a different approach than previous work, measuring resistance to manipulation by whether neural networks of various sizes can learn to profitably manipulate a given voting method in expectation, given different types of information about how other voters will vote. Like the classic results on manipulation, we focus on the case of a single manipulating voter. A single voter can almost never affect the outcome of a large election with thousands of voters or more by changing their vote, so rational manipulation by a single voter is most relevant for small elections in committees, boards, etc. The classic results on manipulation also in effect assume that the manipulator knows exactly how all other voters will vote, which is unrealistic in many voting contexts. By contrast, we train neural networks to manipulate on the basis of different types of limited information.
Related work
Manipulation under Limited Information
@@
A number of previous papers study whether a voter can successfully manipulate in an election under limited information about how other voters will vote (Myerson and Weber 1993; Conitzer, Walsh, and Xia 2011; Reijngoud and Endriss 2012; Meir, Lev, and Rosenschein 2014; Endriss et al. 2016; Lang 2020; Veselova 2023), including with the use of heuristics (Chopra, Pacuit, and Parikh 2004; Laslier 2009; Meir 2018; Fairstein et al. 2019). The information of the manipulator is typically represented by a set of preference profiles, all of which agree on (i) the manipulator’s own preferences and (ii) some other partial information (e.g., all profiles in the manipulator’s information set agree on who will win the election if the manipulator votes sincerely, or all profiles in the set are such that for each of the other voters $i$, $i$’s ranking in the profile extends some known partial order over the candidates that the manipulator attributes to $i$), perhaps supplemented with a probability measure over such profiles (Lu et al. 2012). By contrast, in this paper we will represent limited information by the inputs to a neural network.
Machine Learning and Voting Theory
@@
Several previous works apply machine learning to problems in voting theory, though not in the way we do here. Anil and Bao 2021, Burka et al. 2022, and Matone et al. 2024 study the learnability of various voting methods; Kang, Han, and Xia 2023 studies learning how to explain election results for various voting methods; and Armstrong and Larson 2019 and Firebanks-Quevedo 2020 use machine learning to create new voting methods satisfying desiderata. But none of these papers discuss learning to manipulate as a voter. Manipulation is studied in Airiau, Grandi, and Perotto 2017, but only in the context of iterative voting, [^1] In iterative voting, after all voters submit rankings and a tentative winner is announced, the voters are allowed to sequentially change their rankings, with a new tentative winner announced after each change, until an equilibrium is reached. whereas we focus on learning to manipulate in traditional elections, where the final winner is immediately computed after all voters submit their rankings.
Machine Learnability as a Metric of Task Difficulty
@@
Sufficiently large neural networks are able to learn arbitrarily complex functions, including fitting to random data (Zhang et al. 2017). If the model is not large enough to fully memorize the training data, however, learning the training data requires generalization. In the fields of reinforcement learning and natural language processing, it is commonly held that more complex problems may require larger and more complex networks. Indeed, previous work has shown that model performance grows as the number of learnable parameters increases (Hestness et al. 2017; Golubeva, Neyshabur, and Gur-Ari 2021). In this paper, we use required model size as a proxy for task difficulty. Not only has learnability by neural networks been taken to be suggestive of human learnability (Steinert-Threlkeld and Szymanik 2020), but also humans may use a neural network to help them manipulate an election under limited information.
Preliminaries
@@
Given a set $V$ of voters and a set $X$ of candidates, a preference profile for $(V,X)$ is a function $\mathbf{P}$ assigning to each $i\in V$ a linear order $\mathbf{P}_{i}$ of $X$. Where $\mathbf{Q}$ is a preference profile for $(V,X)$, $i\not\in V$, and $\mathbf{P}_{i}$ is a linear order of $X$, we write $(\mathbf{P}_{i},\mathbf{Q})$ for the preference profile that assigns to $i$ the linear order $\mathbf{P}_{i}$ and assigns to each $j\in V$ the linear order $\mathbf{Q}_{j}$. For a profile $\mathbf{P}$ and voter $i\in V$, let $\mathbf{P}_{-i}$ be the profile for $(V\setminus\{i\},X)$ obtained by restricting $\mathbf{P}$ to all voters except $i$, so we may write $\mathbf{P}=(\mathbf{P}_{i},\mathbf{P}_{-i})$. For a candidate $a\in X$, let $\mathbf{P}_{-a}$ be the profile for $(V,X\setminus\{a\})$ obtained from $\mathbf{P}$ by restricting each voter’s linear order to $X\setminus\{a\}$.
Given a profile $\mathbf{P}$ and candidates $a,b\in X$, the margin of $a$ vs. $b$ in $\mathbf{P}$, denoted $\mathrm{Margin}_{\mathbf{P}}(a,b)$, is the number of voters who rank $a$ above $b$ minus the number who rank $b$ above $a$ in $\mathbf{P}$. A Condorcet winner in $\mathbf{P}$ is a candidate $c\in X$ with a positive margin over every $a\in X\setminus\{c\}$.
@@
A utility profile for $(V,X)$ is a function $\mathbf{U}$ assigning to each $i\in V$ a utility function $\mathbf{U}_{i}:X\to\mathbb{R}$, where we assume that ${\mathbf{U}_{i}(x)\neq\mathbf{U}_{i}(y)}$ whenever $x\neq y$ [^2] The probability models for utility profiles that we use make a tie in the utilities of distinct candidates a measure zero event. Given such a utility profile $\mathbf{U}$, its induced preference profile $\mathbf{P}(\mathbf{U})$ assigns to each $i\in V$ the linear order $\succ_{i}$ defined by
$x\succ_{i}y$ iff $\mathbf{U}_{i}(x)>\mathbf{U}_{i}(y)$. (1)
@@
We say $\mathbf{P}_{i}^{\prime}$ is optimal if the left-hand side of (2) is maximized for $\mathbf{P}^{\prime}_{i}$ among all possible linear orders of $X$. We assume, as in standard decision theory (Kreps 1988), that our manipulating agent aims to maximize expected utility and hence aims to submit an optimal ranking.
A voting method $F$ is manipulable at $\mathbf{U}$ by $i$ if there is some profitable manipulation of $F$ at $\mathbf{U}$ by $i$; and $F$ is manipulable if there is some utility profile $\mathbf{U}$ and voter $i$ such that $F$ is manipulable at $\mathbf{U}$ by $i$. This notion of manipulability of $F$ coincides with the notion of manipulability of $F$ in Gibbard 1977 when we regard $F$ as a probabilistic voting method that assigns to each profile $\mathbf{P}$ the lottery $F_{\ell}(\mathbf{P})$.
Voting Methods
@@
Instant Runoff with parallel-universe tiebreaking (IRV-PUT): if more than half of the voters rank the same candidate $a$ in first place, then $a$ wins; otherwise a candidate $a$ is an IRV-PUT winner if for one of the candidates $b$ who received the fewest first-place votes in $\mathbf{P}$, $a$ is the IRV-PUT winner in the profile $\mathbf{P}_{-b}$.
@@
Black’s: if there is a Condorcet winner, that candidate wins; otherwise the Borda winners win.
Minimax: the winners are those candidates $a$ who minimize the quantity $\mathrm{max}\{\mathrm{Margin}_{\mathbf{P}}(b,a)\mid b\in X\}$.
@@
Plurality, Instant Runoff, and Borda are perhaps the most famous of preferential voting methods. Plurality has been used for many centuries, and Instant Runoff and Borda date back to at least the 18th century. For Instant Runoff, there are multiple ways of handling ties in the number of first-place votes. [^3] Another version of IRV, as in Taylor and Pacelli 2008, simultaneously eliminates all candidates with the fewest first-place votes in a given round, unless all candidates have the same number of first-place votes, in which case all win. The PUT version is popular in computational social choice (Wang et al. 2019).
All of the other methods are Condorcet consistent in the sense that if there is a Condorcet winner, that candidate is the unique winner according to the method. Plurality, Borda, and IRV-PUT all violate Condorcet consistency. Black’s (Black 1958) and Minimax (Simpson 1969; Kramer 1977) are two of the most well known Condorcet methods.
The Nanson [^4] There are two versions of Nanson, one that removes all candidates with below average Borda score (Strict Nanson) and one that removes all candidates with less than or equal to average Borda score (Weak Nanson) unless doing so eliminates everyone. Apparently Nanson himself had in mind Weak Nanson (Niou 1987), but most of the literature in computational social choice focuses on Strict Nanson (see Brandt et al. 2016), so we do as well. method (Nanson 1882) has previously been studied in connection with strategic voting. In Narodytska, Walsh, and Xia 2011, it is shown that the problem of manipulating Nanson (and the related Baldwin method) so as to elect a desired candidate is NP-hard when the number of candidates is allowed to increase.
Finally, we include the recently proposed Split Cycle voting method (Holliday and Pacuit 2023a), whose manipulability has been studied in Durand 2023, as well as one of its refinements, [^5] A voting method $F$ refines a method $G$ if for any profile $\mathbf{P}$, $F(\mathbf{P})\subseteq G(\mathbf{P})$. Other refinements of Split Cycle include Beat Path (Schulze 2011) and Ranked Pairs (Tideman 1987). the Stable Voting method (Holliday and Pacuit 2023b). These methods satisfy not only Condorcet consistency but also the stronger property—violated by Minimax—of Smith consistency, meaning that their winners always belong to the Smith set, the smallest set of candidates such that every candidate inside the set beats every candidate outside the set head-to-head. No previous work has studied the manipulability of Stable Voting, so studying this method tests if our approach can predict the manipulability of a method as measured in other ways in the future.
Learning to Manipulate
@@
We use MLPs as our underlying architecture because they impose no structural assumptions on the input data. This flexibility is particularly important in our experiments, as we compare a variety of types of limited information, each with its own structure (see below). To ensure that our results in this initial investigation are not biased toward any specific type of information, we chose an architecture that does not depend on structural assumptions about sequence and order (as with transformers) or local patterns (as with CNNs).
Implementation Details
We optimize weights $\theta$ of an MLP $f$ whose input $x$ consists of a utility function $\mathbf{U}_{v}$ for our manipulating voter $v$, as well as some limited information $I$ about the full utility profile $\mathbf{U}$ (see below). We apply a softmax to the output of the MLP to generate a probability distribution over all possible actions of $v$, namely the $m!$ possible rankings, labeled 0 through $m!-1$, that $v$ can submit, as in (3):
@@
Note that $v$ is allowed (though there is no special incentive) to submit the sincere ranking given by their utility function.
Probability models for profiles
To generate utility profiles for our experiments described below, we first used a standard uniform utility model (see, e.g., Merrill 1988, p. 16): for each voter independently, the utility of each candidate for that voter is drawn independently from the uniform distribution on the $[0,1]$ interval.
We also used a spatial 2D model: each candidate and each voter is independently placed in $\mathbb{R}^{2}$ according to the multivariate normal distribution (as in Merrill 1988) with no correlation between the two dimensions; the utility of a candidate for a voter is the square of the Euclidean distance between the candidate and the voter (using the quadratic proximity utility function as in Merrill and Grofman 1999, p. 21).
[…]
These utility profiles can then be parameterized as 2D matrices, $U\in\mathbb{R}^{n\times m}$, where $n$ is the number of voters, $m$ is the number of candidates, and $U[i,c]\in\left[0,1\right]$. To select the utility function for a given voter $i$ is to select the row $U_{i}$.
Choices of Limited Information
We experimented with providing different types of input to the MLP, including the following types that are often taken to be natural forms of polling information in voting theory (see, e.g., Reijngoud and Endriss 2012, § 2.2, Veselova 2023):
@@
These additional inputs are flattened and concatenated to $v$’s utility function before being used as input to the MLP.
Note, crucially, that the full preference profile $\mathbf{P}$ is not uniquely determined by any of the types of limited information above. If the manipulating voter $v$ had full knowledge of $\mathbf{P}$, they could simply compute which of the $m!$ linear orders would be optimal to submit given $\mathbf{P}_{-i}$. But if $v$ has only some limited information $I$, e.g., the margin matrix, then there is a set $\mathbb{P}(I)$ of preference profiles that could have generated the given limited information $I$. Where $\mu$ is the probability distribution on preference profiles induced by our probability distribution on utility profiles, it might not be feasible for $v$ to sample sufficiently many profiles from the conditional distribution $\mu(\cdot\mid\mathbb{P}(I))$ in order to obtain a good estimate of the expected utility of submitting a linear order, when the limited information is $I$, for each of the $m!$ linear orders. [^6] For some types of limited information, such as the plurality ranking, a simple rejection sampling approach is feasible: sample profiles according to $\mu$ and throw out those that do not realize the given plurality ranking. Since there are relatively few plurality rankings, one rapidly acquires many profiles realizing a given plurality ranking. By contrast, since there are vastly more margin matrices, it may be necessary to sample a huge number of profiles before finding sufficiently many that realize a given margin matrix. Rather than investigating such a sampling approach, in this paper we will train an MLP to input $I$ and output an optimal ranking given $v$’s utilities.
@@
  • optimizing labeling: all optimal rankings to submit are labeled by $1$, and all other rankings are labeled by 00.777We duplicated our experiments with a satisficing labeling (if there are profitable manipulations, they are labeled by $1$, and all others rankings are labeled by 00; otherwise all rankings that do at least as well as the sincere ranking are labeled by $1$ and all others by 00), but MLPs trained with this labeling had qualitatively indistinguishable results from those trained with the optimizing labeling.
The output of our MLPs is a distribution over all $m!$ rankings given some information $x$ about the current utility profile. It is equally valid for our agent to choose any of the positively-labeled rankings. We treat the binary labelings as a mask over the rankings and reduce the distribution $\pi(x)$ to two values: the probability of choosing a positively-labeled ranking or not. We compute the final loss as the mean-squared error between this reduced distribution and the distribution assigning probability 1 to choosing a positively-labeled ranking and 0 to choosing a non-positively-labeled ranking. [^8] After a first run of our experiments, we reran the entire experiment for the uniform utility model with the final loss computed as the binary cross-entropy loss, but this did not lead to any substantial difference in the performance of trained MLPs that would affect our conclusions (see the GitHub repository for details).
Evaluation
@@
  • argmax: select the ranking with the maximum probability in the output of the MLP.
@@
The normalization in (4) is the standard normalization for relative utilitarianism (Dhillon and Mertens 1999), which we use to compare utility differences across profiles. Indeed, (4) is equivalent to taking the difference in the Kaplan-normalized (d’Aspremont and Gevers 2002, p. 470) expected utilities of submitting $\mathbf{P}^{\prime}$ and of submitting $\mathbf{P}$, where $\hat{u}=\mathrm{max}(\{\mathbf{U}_{i}(x)\mid x\in X\})$, $\check{u}=\mathrm{min}(\{\mathbf{U}_{i}(x)\mid x\in X\})$:
@@
For a given trained MLP, we sample utility profiles according to one of our probability models and compute the average profitability of the MLP’s submitted rankings. We take the average profitability that is achieved against a given voting method as a measure of a voter’s incentive to manipulate against that voting method. This is a more revealing measure than the frequency with which the trained MLP manipulates against a voting method, since many of the MLP’s manipulations may be no better than sincere voting.
@@
For the number of profiles to sample, we continued sampling until the estimated standard error of the mean (Robert and Casella 2004, § 3.2) for profitability fell below 5e-4, resulting in small error bars (see Figure 2). [^9] The number of profiles sampled for evaluation of an MLP was at least 4,096 with an average of approximately 36,000 and a standard deviation of approximately 31,000.
Baselines